OpenStudy (anonymous):

Solve by using the quadratic formula to solve x2 + 6x = –3.

OpenStudy (anonymous):

\[x^2+6x+3=0\]\[\Large x_{1/2}\frac{-6 \pm \sqrt{36-12}}{2} =\frac{-6\pm \sqrt{24}}{2}\] now continue...

OpenStudy (anonymous):

is it